# Steering AI-Driven Personalization of Scientific Text for General Audiences

### v1

@@ 6.4. Limitations and Future Studies
Our study had several limitations. First, it was conducted in a controlled setting. Specifically, a small sample size of only 15 participants were asked to explore two pre-selected science articles from health and environment domains using our tool. Future studies should consider a large-scale in-the-wild study where participants could choose from a wider range of scientific articles. Secondly, our study setup required participants to provide feedback based on brief interactions with the tool. The long-term effects of interactive reading experiences with AI-personalized translation tools may differ from our findings. Future longitudinal studies are warranted to provide a more comprehensive understanding of how users interact with and benefit from these types of tools over extended periods. Lastly, our participant demographics were limited, with most participants (12 out of 15) being in their twenties and recent college graduates. Only three participants were from other age groups. This age-homogeneous participant pool may have influenced the findings of our study. Future studies should include participants from diverse backgrounds in terms of age, education level, and culture to capture a broader range of perspectives on AI-personalized translation tools. For instance, the impact of AI-personalized translation tools in high school science learning classes (Lee and Perret, 2022) or children-parent learning contexts (Cingel and Piper, 2017) could be different from our findings.
@@ 7. Conclusion
Our study investigated the potential of AI-personalized translations to enhance science communication in digital media platforms. We developed TranSlider, an AI-powered reading interface that generates personalized translations of scientific text for general audiences. The tool’s key feature involves an interactive slider that allows users to adjust the degree of personalization to generate tailored translations. Through an exploratory study with 15 participants, we found that participants preferred varying degrees of personalization for the tailored translations, from deeply relatable and contextualized translations to short and slightly contextualized. We also found that reading multiple translations enabled them to have a compounding understanding of scientific content. We present several implications in supporting communication across diverse backgrounds in collaborative contexts and highlight the importance of user agency in AI-powered systems to facilitate desirable outputs.

### v2

@@ 7.3. Limitations and Future Work
Our study had several limitations. First, it was conducted in a small controlled setting. Specifically, a pool of 15 participants were asked to explore two pre-selected science articles from health and environment domains using our tool. Future studies should consider a large-scale in-the-wild study where participants could choose from a wider range of scientific articles. Secondly, our participants might not fully represent the general audience with varying degrees of formal education. More research is needed by studying participants with diverse educational levels (e.g.,  high school graduates or those with lower educational attainment), to better understand how people with lesser exposure to formal education could draw benefit from our tool. Similarly, our participant demographics were limited, with most participants (12/15) in their twenties and recent college graduates. Only three participants were from other age groups. This age-homogeneous participant pool may have influenced the findings of our study. Future studies should include participants from diverse backgrounds in terms of age, education level, and culture to capture a broader range of perspectives on AI-personalized translation tools. For instance, the impact of AI-personalized translation tools in high school science learning classes (Lee and Perret, 2022) or children-parent learning contexts (Cingel and Piper, 2017) could be different from our findings. Lastly, our study setup required participants to provide feedback based on brief interactions with the tool. The long-term effects of interactive reading experiences with AI-personalized translation tools may differ from our findings. Future longitudinal studies are warranted to provide a more comprehensive understanding of how users interact with and benefit from these types of tools over extended periods.
@@ 8. Conclusion
Our study investigated the potential of AI-personalized translations to enhance science communication in digital media platforms. We developed TranSlider, an AI-powered reading interface that generates personalized translations of scientific text for general audiences. The tool’s key feature involves an interactive slider that allows users to adjust the degree of personalization to generate tailored translations. Through an exploratory study with 15 participants, we found that participants preferred varying degrees of personalization for the tailored translations, from deeply relatable and contextualized translations to short and slightly contextualized ones. Interestingly, reading multiple translations enabled them to have a compounding understanding of scientific content. We present several implications for science communication and for supporting communication across individuals with diverse backgrounds in collaborative contexts. We also discuss designing flexible interfaces that allow users to preserve control over AI tools to steer desired outcomes. We hope our study opens a new avenue to promote science communication for general audiences.
Acknowledgements.
We thank all reviewers for their constructive feedback. We also thank Accenture Labs’ leadership for supporting this research and Labs’ Digital Experiences team members for continued guidance. Furthermore, special thanks to members from the broader research community, including Elizabeth Churchill, Jung Wook Park, John Chung, Duri Long, and Tal August for their insightful discussion. We also greatly appreciate Matt Kay for his thoughtful comments and suggestions on the manuscript and Jessica Hullman for introducing extensive related work in this domain.
Appendix A Original Science Articles
A.1. Health (Kong et al., 2024)
Title: A glycolytic metabolite bypasses ‘‘two-hit’’ tumor suppression by BRCA2
Abstract: Knudson’s ‘‘two-hit’’ paradigm posits that carcinogenesis requires inactivation of both copies of an autosomal tumor suppressor gene. Here, we report that the glycolytic metabolite methylglyoxal (MGO) transiently bypasses Knudson’s paradigm by inactivating the breast cancer suppressor protein BRCA2 to elicit a cancer-associated, mutational single-base substitution (SBS) signature in nonmalignant mammary cells or patient-derived organoids. Germline monoallelic BRCA2 mutations predispose to these changes. An analogous SBS signature, again without biallelic BRCA2 inactivation, accompanies MGO accumulation and DNA damage in Kras-driven, Brca2-mutant murine pancreatic cancers and human breast cancers. MGO triggers BRCA2 proteolysis, temporarily disabling BRCA2’s tumor suppressive functions in DNA repair and replication, causing functional haploinsufficiency. Intermittent MGO exposure incites episodic SBS mutations without permanent BRCA2 inactivation. Thus, a metabolic mechanism wherein MGO-induced BRCA2 haploinsufficiency transiently bypasses Knudson’s two-hit requirement could link glycolysis activation by oncogenes, metabolic disorders, or dietary challenges to mutational signatures implicated in cancer evolution.
A.2. Environment (Guelfo et al., 2024)
Title: Lithium-ion battery components are at the nexus of sustainable energy and environmental release of per- and polyfluoroalkyl substances
Abstract: Lithium-ion batteries (LiBs) are used globally as a key component of clean and sustainable energy infrastructure, and emerging LiB technologies have incorporated a class of per- and polyfluoroalkyl substances (PFAS) known as bis-perfluoroalkyl sulfonimides (bis-FASIs). PFAS are recognized internationally as recalcitrant contaminants, a subset of which are known to be mobile and toxic, but little is known about environmental impacts of bis-FASIs released during LiB manufacture, use, and disposal. Here we demonstrate that environmental concentrations proximal to manufacturers, ecotoxicity, and treatability of bis-FASIs are comparable to PFAS such as perfluorooctanoic acid that are now prohibited and highly regulated worldwide, and we confirm the clean energy sector as an unrecognized and potentially growing source of international PFAS release. Results underscore that environmental impacts of clean energy infrastructure merit scrutiny to ensure that reduced CO2 emissions are not achieved at the expense of increasing global releases of persistent organic pollutants.